For the diprotic weak acid H2A, Ka-3.6x10 and Ka2 8.6 x10 What is the pH of a 0.0700 M solution of H2A? What are the equilibrium concentrations of H2A and A in this solution? Number pH- Number H,A Number
Answer #1

-6 2. 0100- [½m] = o-o700-r o.oe7oo._ o.coo51 | cse.nom -log (o.ooot) = 3-3 一.log [H+3 ka = 15-2